# Can Test-time Computation Mitigate Memorization Bias in Neural Symbolic Regression?

### v1

@@ E.1.2 Log-precision Transformers
Definition 2 ($(D,d)$-bounded log-precision Transformer).
@@
For specific definitions of operations that enable approximation in $O(\log k)$ bits, please refer to Section $4$ and Appendix A of Merrill and Sabharwal [19]. We introduce the simulation guarantees for bounded-depth log-precision Transformers as follows.
Lemma 1 (Circuit simulation [19, Cor. 2.1]).
@@
E.1.3 The Boolean Formula Value Problem
Following the definition by Buss [6], we introduce the definition of the Boolean formula value problem as follows.
Definition 3 (Boolean formula value problem).
@@ E.1.3 The Boolean Formula Value Problem
    If $\mathbf{t}_{1}$ and $\mathbf{t}_{2}$ are two Boolean formulae, then ($(\lnot\mathbf{t}_{1}),(\mathbf{t}_{1}\land\mathbf{t}_{2}),(\mathbf{t}_{1}\lor\mathbf{t}_{2})$ are also Boolean formulae.
When given a boolean formula $\mathbf{t}$, the goal of the Boolean formula value problem is to compute whether the evaluation result $\operatorname{eval}(\mathbf{t})$ of a given Boolean formula is 00 or $1$.
E.2 Main Theorem
Prior to proving the main theorem, we state the following Lemma from Feng et al. [8]. The detailed proof of this Lemma can be found in the same paper. The Lemma states that $\mathsf{TC}^{0}$ circuits are capable of identifying the indexes of paired brackets in a string.
Lemma 2 (Bracket parsing [8, Lem. D.3]).
@@ E.2 Main Theorem
$\boldsymbol{g}_{i}(\mathbf{t})=\begin{cases}(-1,j)&\text{if $t_{i}$ is a left bracket and $t_{i},t_{j}$ are paired.}\\ (j,-1)&\text{if $t_{i}$ is a right bracket and $t_{i},t_{j}$ are paired.}\\ (j,k)&\text{if $t_{i}$ is not a bracket, and $t_{j},t_{k}$ are the nearest paired brackets containing $t_{i}$.}\end{cases}$
@@
Theorem 2 (Bounded log-precision Transformer lower bound).
@@
By Lemma 1, $\mathsf{TF}_{\theta}$ can be simulated by a $\mathsf{TC}^{0}$ circuit family of size $\operatorname{poly}(m)$. Hence, under our assumption, $\operatorname{LastTokenPrediction}(m)\in\mathsf{TC}^{0}$.
@@
      replace 00 with {×$\{\times,x_{1},x_{2}\}$ and $1$ with $\{-,x_{1},x_{2}\}$;
@@
      prepend $+$ to $\mathbf{s}$ to form the incomplete token sequence $\widetilde{\mathbf{s}}$;
@@
[…]
@@
$\begin{dcases}\mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},x_{1})}(\mathbf{x}))=\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+x_{1,i}))^{2}=\dfrac{1}{2}((-1)^{2}+(-1)^{2})=1,\\ \mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},x_{2})}(\mathbf{x}))=\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+x_{2,i}))^{2}=\dfrac{1}{2}(0^{2}+0^{2})=0,\\ \mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},C)}(\mathbf{x}))=\underset{c\in\mathcal{C}}{\operatorname{argmin}}~{}\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+c))^{2}=\underset{c\in\mathcal{C}}{\operatorname{argmin}}~{}\dfrac{1}{2}((-c)^{2}+(-1-c))^{2})\geq\dfrac{1}{4},\end{dcases}$
@@
The Boolean formula value problem is $\mathsf{NC}^{1}$-complete under $\mathsf{AC}^{0}$ reductions [6, Thm. 9]. Hence Step 2 and Step 3 indicate that $\operatorname{LastTokenPrediction}(m)\notin\mathsf{TC}^{0}$, contradicting Step 1 and the assumed strict inclusion $\mathsf{TC}^{0}\subsetneq\mathsf{NC}^{1}$. Therefore, such a Transformer cannot exist. ∎
E.3 PAC approximation via iterated self-verification
@@
    Hypothesis class. Fix a maximum depth $D_{0}$ and a grid spaced in $\varepsilon/2$ on $[-1,1]$.
@@
Then, Under Assumption 1, for any $\varepsilon,\delta\in(0,1)$,
@@
(i) Burn-in. There are $K\!\leq\!2^{D_{0}}$ distinct sub-trees of $e^{\star}$. Drawing $r\geq 1$ uniform sub-trees per round, the probability a fixed sub-tree is never drawn in $B$ rounds is $(1-\tfrac{r}{U})^{B}\leq e^{-rB/U}\leq\delta/(2K)$. A union bound over all $K$ sub-trees implies that, after $B$ rounds, the prompt contains every sub-tree of $e^{\star}$ with probability at least $21-\delta/$ 2.
@@
Pr[miss in all $\Pr[\text{miss in all }R\text{ rounds}]\;\leq\;(1-\beta)^{R}\;\leq\;e^{-\beta R}\;\leq\;\delta/2,$
@@
Appendix F Additional Related Work
The GP framework is a traditional and widely used framework for solving symbolic regression. The GP algorithm is a method based on evolutionary computation; initially, several mathematical expressions are formed randomly, and subsequently the expressions are “evolved” by operations such as recombining two expressions, mutating an expression, and eliminating inappropriate expressions [5, 22, 34, 7].
An example of using brute-force algorithms for symbolic regression is AI Feynman [28, 29]. In AI Feynman, neural networks were used to identify properties such as symmetry and separability within given numerical data. These properties were then used to recursively simplify the problem, ultimately reducing it to a form amenable to brute-force solutions.
Petersen et al. [21] proposed Deep Symbolic Regression (DSR), a method that uses reinforcement learning to tackle symbolic regression. In this approach, the authors used a recurrent neural network (RNN) to generate equations as token sequences, with the parameters that govern the selection of the token learned through reinforcement learning. Studies such as Symbolic Physics Learner [25] and Reinforcement Symbolic Regression Machine [36] also use reinforcement learning, where Monte Carlo Tree Search (MCTS) is applied to discover expressions.
Some studies combine several approaches for symbolic regression. For example, neural-guided genetic programming [20] integrates DSR and genetic programming (GP), while the Unified DSR Framework [14] combines GP, AI Feynman, DSR, linear models, and NSR.

### v2

@@ G.1.2 Log-precision Transformers
Definition G.1 ($(D,d)$-bounded log-precision Transformer).
@@
For specific definitions of operations that enable approximation in $O(\log k)$ bits, please refer to Section $4$ and Appendix A of Merrill and Sabharwal (2023b). We introduce the simulation guarantees for bounded-depth log-precision Transformers as follows.
Lemma G.2 (Circuit simulation (Merrill and Sabharwal, 2023b, Cor. 2.1)).
@@
G.1.3 The Boolean Formula Value Problem
Following the definition by Buss (1987), we introduce the definition of the Boolean formula value problem as follows.
Definition G.3 (Boolean formula value problem).
@@ G.1.3 The Boolean Formula Value Problem
    If $\mathbf{t}_{1}$ and $\mathbf{t}_{2}$ are two Boolean formulae, then $(\lnot\mathbf{t}_{1}),(\mathbf{t}_{1}\land\mathbf{t}_{2}),(\mathbf{t}_{1}\lor\mathbf{t}_{2})$ are also Boolean formulae.
When given a boolean formula $\mathbf{t}$, the goal of the Boolean formula value problem is to compute whether the evaluation result $\operatorname{eval}(\mathbf{t})$ of a given Boolean formula is 0 or $1$.
G.2 Main Theorem
Prior to proving the main theorem, we state the following Lemma from Feng et al. (2023). The detailed proof of this Lemma can be found in the same paper. The Lemma states that $\mathsf{TC}^{0}$ circuits are capable of identifying the indexes of paired brackets in a string.
Lemma G.4 (Bracket parsing (Feng et al., 2023, Lem. D.3)).
@@ G.2 Main Theorem
[…]
@@
Theorem G.5 (Bounded log-precision Transformer lower bound).
@@
By Lemma G.2, $\mathsf{TF}_{\theta}$ can be simulated by a $\mathsf{TC}^{0}$ circuit family of size $\operatorname{poly}(m)$. Hence, under our assumption, $\operatorname{LastTokenPrediction}(m)\in\mathsf{TC}^{0}$.
@@
      replace 0 with $\{\times,x_{1},x_{2}\}$ and $1$ with $\{-,x_{1},x_{2}\}$;
@@
      prepend ++ to $\mathbf{s}$ to form the incomplete token sequence $\widetilde{\mathbf{s}}$;
@@
To perform the first step, for all $\lor$ in $\mathbf{t}$, we must replace the nearest left bracket containing $\lor$ with $\lnot(\lnot$ and also replace $\lor$ with $\land\lnot$. By using the results of Lemma G.4, it follows that this operation can be performed by a circuit within $\mathsf{TC}^{0}$ complexity. The second step can be implemented by $\mathsf{AC}^{0}$ circuits, according to Buss (1987, Cor. 11). Since the third and fourth steps only involve replacing and extending obtained sequences, these steps can also be implemented by $\mathsf{AC}^{0}$ circuits.
@@
$\begin{dcases}\mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},x_{1})}(\mathbf{x}))=\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+x_{1,i}))^{2}=\dfrac{1}{2}((-1)^{2}+(-1)^{2})=1,\\ \mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},x_{2})}(\mathbf{x}))=\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+x_{2,i}))^{2}=\dfrac{1}{2}(0^{2}+0^{2})=0,\\ \mathcal{L}(\mathbf{y},e_{(\widetilde{\mathbf{s}},C)}(\mathbf{x}))=\underset{c\in\mathcal{C}}{\operatorname{argmin}}~\dfrac{1}{2}\sum_{i=1}^{2}(y_{i}-(1+c))^{2}=\underset{c\in\mathcal{C}}{\operatorname{argmin}}~\dfrac{1}{2}((-c)^{2}+(-1-c)^{2})\geq\dfrac{1}{4},\end{dcases}$
@@
The Boolean formula value problem is $\mathsf{NC}^{1}$-complete under $\mathsf{AC}^{0}$ reductions (Buss, 1987, Thm. 9). Hence Step 2 and Step 3 indicate that $\operatorname{LastTokenPrediction}(m)\notin\mathsf{TC}^{0}$, contradicting Step 1 and the assumed strict inclusion $\mathsf{TC}^{0}\subsetneq\mathsf{NC}^{1}$. Therefore, such a Transformer cannot exist. ∎
G.3 PAC approximation via iterated self-verification
@@
    Hypothesis class. Fix a maximum depth $D_{0}$ and a grid spaced in $\varepsilon/2$ on [1,1][-1,1].
@@
Then, Under Assumption G.6, for any $\varepsilon,\delta\in(0,1)$,
@@
(i) Burn-in. There are $K\!\leq\!2^{D_{0}}$ distinct sub-trees of $e^{\star}$. Drawing $r\geq 1$ uniform sub-trees per round, the probability a fixed sub-tree is never drawn in $B$ rounds is $(1-\tfrac{r}{U})^{B}\leq e^{-rB/U}\leq\delta/(2K)$. A union bound over all $K$ sub-trees implies that, after $B$ rounds, the prompt contains every sub-tree of $e^{\star}$ with probability at least $1-\delta/2$.
@@
$\Pr[\text{miss in all }R\text{ rounds}]\;\leq\;(1-\beta)^{R}\;\leq\;e^{-\beta R}\;\leq\;\delta/2,$